If at 1.00 atm of pressure water boils at 100°C, at what temperature would water boil if the pressure is 600 torr? (This shows why food doesn't cook well at higher elevations) (Remember Torr is the same as mmHg)


Law: ______________________


Answer:____________________


If the winter pressure that most car tire tires should be at to wear evenly is 32.0 psi, what is the pressure in atm?


Law: Not a Las, just a simple Conversion


Answer:____________________



A gas is placed in a balloon with a volume of 3.0 L at 28 °C and 900 mmHg. What would be the new volume for the gas if placed under STP? (complete after Friday’s lesson)


Law: ______________________


Answer:____________________


1
Expert's answer
2021-04-28T02:46:13-0400

1) Gay-Lussac’s Law

1torr = 0.00132 atm

= 600 × 0.00132

= 0.792 atm

P1/T1 = P2/T2

1/373 = 0.792/Y

0.002681Y = 0.792

= 295.41K

295.41-273

= 22.41°C

2.Pressure conversion

1psi = 0.068atm

= 0.068 × 32

= 2.176 atm

3. Combined gas laws

V2 = P1V1T2/P2T1

= 900 × 273 × 3/301 × 760

= 3.2 L